A producer of felt-tip pens has received a forecast of demand of 33,000 pens for the coming month from its marketing department. Fixed costs of $27,000 per month are allocated to the felt-tip operation, and variable costs are 25 cents per pen.

a. Find the break-even quantity if pens sell for $2 each. (Round your answer to the next whole number.) QBEP 15,428 15,428 Correct units

b. At what price must pens be sold to obtain a monthly profit of $11,000, assuming that estimated demand materializes? (Round your answer to 2 decimal places. Omit the "$" sign in your response.)

Price $
Answer #1

Answer a: 15,428

Explanation: Demand= 33000 pens

Fixed operation cost= 27000

variable cost= 0.25 per pen

Selling price of pen= 2

Profit per pen= Selling price-variable cost= 2-0.25= $1.75

the breakeven quantity= Fixed cost/profit per pen= 27000/1.75= 15,428.57

Answer b: $1.40 per pen

Explanation: Total cost of 33000 pens= fixed cost+ (variable cost*33000) = 27000+ (0.25*33000)= $35,250

Required profit= $11000,

Therefore total sales required= 35250+ 11000= $46,250

Selling price required = total sales/33000= 46250/33000= $1.40 per pen